# MCQG-SRefine: Multiple Choice Question Generation and Evaluation with Iterative Self-Critique, Correction, and Comparison Feedback

### v2

@@ 2.4 Critique
After generating all components in INIT and QA step, the LLM is asked to critique each component. The set to be critiqued is $u_{\text{critique}}=\langle c,q,a,d,r\rangle$. The LLM receives a scoring guide $G$, which includes all aspects that need to be evaluated for each component. The prompt includes this guide $G$ as well as several manually written example critiques of scored components $E^{\text{fs}}_{\text{critique}}=\langle e^{\text{fs}}_{c},e^{\text{fs}}_{q},e^{\text{fs}}_{a},e^{\text{fs}}_{r}\rangle$ and $\langle G,n,t,k,c,q,a,d,a_{a}\rangle$. The final output of this step is LLM critique feedback on all components, $f=\langle f_{c},f_{q},f_{a},f_{d},f_{r}\rangle$, which includes short text critiques and scores for each aspect. The aspects for scoring different components in $G$ are as follows: Context: Relevant, Concise, Coherent, Consistent, Specific, Fluent, Clueing, Completeness, and Misdirection. Question: Relevant, Clear, Concluding, Difficulty, and Clarity. Correct Answer: Relevant, Occurrence, Justification, Depth of Understanding, and Prevention of Guesswork. Distractors: Format, Length, Relation, Variation, Plausibility, Differentiation, and Common Mistakes. Chain of Thought/Reasoning: Logical Flow, Evidence-Based Reasoning, and Consideration of Options. We provide detailed explanations for each aspect of every component in Appendix Table LABEL:tab:demonstration_MCQG-SRefine, and LLM-Critique prompts in Table LABEL:prompt. The total score for each component is calculated by summing up all individual aspect scores. A sample output is provided in Appendix Table LABEL:tab:demonstration_MCQG-SRefine.
@@ 3 Experimental Design and Setup
For the medical cases utilized in the generation of USMLE questions, we used unidentified patient summaries from the PMC-Patients dataset Zhao et al. (2023). The average length of these patient summaries was ~419 words. The frequency of topics used is listed in Appendix Table 13 14.
@@
To assess the quality of the questions (RQ1), we first engaged two medical experts  [^8] Two medical students with 2+ years hospital experience. to express their preferences between the two sets (GPT-4 and MCQG-SRefine) of system-generated questions based on an annotation guideline (Appendix Table LABEL:tab:human_eval_guide). The evaluators were blinded to the source of the questions, and the order of the questions was randomized for each data point. We calculated the Percentage Agreement (87.5%) and Cohen’s kappa (0.66722) for the two evaluators’ preferences, indicating substantial reliability of our human evaluation settings. Subsequently, a third human expert  [^9] One licensed physician. facilitated discussions with the initial evaluators to make the final decision for each data point, representing the final human expert preference (referred to as Expert X).
@@ 4 Results
Figure 3: The quality expert preference for the GPT-4 and the MCQG-SRefine question. The data is divided into Human and Machine based on how the topic $t$ and key points $k$ were generated. We only put the final Expert X preferences here, but we provide more results in the Appendix Table 10. The percentage agreement between experts is 87.5% (Human<$t$, $k$>: 90%, Machine<$t$, $k$>: 85%). The Cohen’s kappa between experts is 0.66722 (Human<$t$, $k$>: 0.75, Machine<$t$, $k$>: 0.57), indicating substantial reliability.
@@
Figure 3 demonstrates the overwhelming advantage of MCQG-SRefine over GPT-4 with a 70-80% win rate in human preference about question quality (RQ1). In Figure 4, we also observe a decrease of about 80% in easy questions, a 2.25 times increase in medium questions, and a 4 times increase in hard questions with the input of these medical cases with expert-provided topics and key points. For machine-provided topics and key point cases, the proportion of easy questions decreased by 33.3%, a 2 times increase in medium questions, but there was no increase in the proportion of hard questions (RQ2). This demonstrates the effectiveness of MCQG-SRefine in increasing the difficulty of questions. This also indicates that the quality of topics and key points provided by experts is higher, so LLM can generate more difficult questions by thinking more deeply during the critique and correction steps. Further improving the quality of the machine-provided topics and key points can be a future direction for improvement.
@@
Figure 5 and Table 1 present the results of the LLM-as-judge evaluation. For human-provided topics and key point cases, MCQG-SRefine achieved a win rate of 79.8% compared to 20.2% for GPT-4. Similarly, for machine-provided topics and key point cases, MCQG-SRefine achieved a win rate of 80.1%, outperforming GPT-4’s 19.9%. Notably, MCQG-SRefine consistently outperformed GPT-4 across all five evaluated components with LLM-as-judge (Rating) results rather than demonstrating an advantage in only a single aspect  [^12] The aspect-level score can be found in Appendix Table 11.. This indicates a balanced and comprehensive improvement of MCQG-SRefine.
Figure 4: The difficulty expert evaluation for the GPT-4 generated question and the MCQG-SRefine question.
@@
Qualitative analysis One of the main issues with the questions generated by GPT-4 is that they often directly include the correct answer or too obvious relevant keywords within the context component (Appendix Table LABEL:tab:human_eval_case Case Studies 1 and 3). So the questions directly generated by GPT-4 often make the answers obvious, but MCQG-SRefine can modify this information into hints for the correct answer through its critical and corrective steps, which experts consider a better way to construct questions for candidates (Case Study 4).
@@
Another finding is that the questions generated by MCQG-SRefine are more concise compared to those by GPT-4 (Case Study 1). Experts pointed out that this conciseness makes them more similar to real USMLE questions. Our experiments show that GPT-4 adopted a very conservative strategy when generating context due to our emphasis on hallucination issues in the prompts. This strategy involves copying and pasting much information from raw medical case inputs to avoid generating potentially incorrect new facts. Although this does reduce the occurrence of hallucinations - our human assessment shows that 7.5% of GPT-4 problems contain factual errors - it inevitably sacrifices the typical simplicity and highly refined information presentation of USMLE problems, as well as the logical coherence of multi-hop reasoning between information in context, question, correct answer, and distractors. This is a significant stylistic difference between USMLE questions and the original medical cases. We also found that prompt engineering was ineffective in resolving this issue. We interpret this as a shortcut behavior learned by GPT-4 during aligning with human preferences stage training (e.g., RLHF Ouyang et al. (2022)) to reduce output diversity to mitigate hallucinations Kirk et al. (2023). In contrast, MCQG-SRefine maintains a high level of factual accuracy (with 5% factual errors) and further improves the quality of USMLE questions by iteratively criticizing and correcting each component of the question, as well as shifting the perspective from QG to QA to make it closer to actual exam questions. This is another major reason why MCQG-SRefine outperforms GPT-4 in human evaluation. However, when MCQG-SRefine reduced the amount of useless information provided in the questions to increase quality and difficulty, omitting too much information sometimes made inferring the correct answer more challenging (Case Study 2), although this was a rare occurrence in our human evaluation (7.5%).
@@
Table 1: The win rate of MCQG-SRefine and GPT-4 questions using LLM-as-Judge (Comparision).
5 Discussion
@@ 6.1 LLMs for Generating Medical MCQs
Cheung et al. (2023) conducted the first comprehensive study comparing LLMs with human experts for generating medical exam MCQs. Using the ChatGPT-3.5 interface, they generated MCQs from two standard medical textbooks Kasper et al. (2015); Williams and O’Connell (2008). Each question included four answer options, without any post-generation modifications. ChatGPT took only 21 minutes to generate 50 MCQs, roughly 10% of the time required by human experts (211 minutes). Despite the efficiency, human-authored questions were found to be superior in quality, outperforming AI in 60% of the acceptable MCQs.
Klang et al. (2023) performed a blind evaluation of MCQs generated by ChatGPT-4. Initially, ChatGPT-4 generated questions with four answer options using examples from previous exams (few-shot learning). These questions were brief and lacked clinical context, necessitating an additional prompt to include clinical background. After refinement, most of the MCQs were considered effective.
Agarwal et al. (2023) compared ChatGPT-3.5, Bard, and Bing in generating MCQs for an 11-module physiology course by the National Medical Commission (NMC) of India. The results indicated that ChatGPT generated more effective MCQs, although they were less challenging compared to those generated by Bard and Bing.
Ayub et al. (2023) focused on dermatology board exams by uploading Continuing Medical Education (CME) articles from the Journal of the American Academy of Dermatology (JAAD) into ChatGPT-3.5. The generated MCQs were evaluated based on accuracy, complexity, and clarity, with only 40% meeting the criteria for applicability, while the remainder were found to be unclear, inaccurate, or overly simplistic.
Sevgi et al. (2023) tasked ChatGPT-3.5 with generating three neurosurgery board exam questions, including answers and explanations. Similarly, Han et al. (2024) used ChatGPT-3.5 to generate three MCQs incorporating clinical context and lab values, experimenting with rephrasing to adjust answers and increase difficulty.
Totlis et al. (2023) employed ChatGPT-4 to generate MCQs on anatomy, focusing on increasing difficulty levels and creating appropriate answer pairings. Biswas (2023) used ChatGPT to prepare USMLE Step 1 MCQs, though no independent evaluations were conducted for these studies.
To our knowledge, this study is the first to comprehensively explore both the qualitative and quantitative aspects of converting real-world medical cases into USMLE-style MCQs. We leverage authentic medical case information to generate high-quality, up-to-date questions that align with current medical standards and best practices.
@@ 6.2 Self-Refine using LLM Feedback
Learning from feedback helps align LLMs with desired outcomes, enhancing their ability to follow instructions through different forms of feedback, such as human preference feedback Ouyang et al. (2022), AI-generated preference feedback Lee et al. (2023); Dubois et al. (2024), or fine-grained feedback Wu et al. (2024); Lightman et al. (2023).
Unlike preference and fine-grained feedback, which provide scalar values as training signals, natural language or correction feedback offers richer information Scheurer et al. (2022); Ma et al. (2023); Yao et al. (2023); Mishra et al. (2024), making it particularly effective for self-correcting language models Welleck et al. (2022); Pan et al. (2023).
Recent research has demonstrated that LLMs Achiam et al. (2023); Heng et al. (2024) can self-correct their responses to meet various user requirements, such as reducing harmful content, incorporating specific keywords, diversity requirement generation, or debugging code Madaan et al. (2024); Chen et al. (2023b). This self-correction process generally involves generating a critique that identifies shortcomings in the initial response, followed by revising it based on the self-critique—an iterative process that has shown promise in enhancing LLM output quality Pan et al. (2023).
Inspired by the success of these iterative self-refinement methods, we are the first to explore using this approach for generating USMLE-style MCQs. By leveraging self-feedback, we aim to create high-quality, clinically relevant questions that adhere to the rigorous standards of medical education.
@@ 6.3 LLM-as-Judge using LLM feedback
The critique capabilities of LLMs have been extensively used for the automatic evaluation of response quality, often employing models like GPT-4 Achiam et al. (2023); Liu et al. (2023); Fu et al. (2023) or critique-adjusted LLMs Ke et al. (2023); Li et al. (2023). Despite their success, these methods have demonstrated instability in certain complex task scenarios Wang et al. (2023b); Zhang et al. (2024b).
LLMs have shown a high correlation with human evaluations in tasks such as summarization and story generation, effectively scoring candidate texts or comparing them based on specific evaluation aspects Chen et al. (2023a). For example, studies by Chiang and Lee (2023) and Kocmi and Federmann (2023) have shown that LLM evaluations produce results comparable to those of expert human evaluators in story generation and translation tasks. Similarly, research by Zheng et al. (2024) and Zhang et al. (2024a) indicates that powerful LLM reviewers, such as GPT-4, achieve over 80% consistency with human preferences in multi-turn dialogue scenarios, both in controlled and crowdsourced settings, reaching agreement levels similar to human evaluators.
Further evidence of GPT-4’s effectiveness as an evaluator is demonstrated by its performance in the PROMETHEUS Kim et al. (2023, 2024) and CRITICBENCH Lan et al. (2024) benchmarks. However, most LLM-as-judge research has primarily focused on general NLP fields, with limited exploration in specialized domains like clinical NLP. This gap is largely due to challenges such as the need for domain-specific knowledge, difficulties in designing evaluation prompts that meet domain standards, and the inherent challenges of generalizing evaluation tools to specialized fields Yao et al. (2024); Singhal et al. (2023); Li et al. (2024).
To our knowledge, we are the first to explore the application of LLM-as-judge in the context of USMLE-MCQ generation and evaluation. Our work addresses these challenges to create reliable, domain-specific evaluations that meet the unique requirements of medical education.
@@ 8 Limitations and Ethical Considerations
Our research focuses on generating medical exam questions, with promising results, though its adaptability to other domains is yet to be explored. We evaluated the method with three medical experts, but future involvement of specialized USMLE exam experts could enhance credibility. Future research should address privacy, fairness, generalizability to other languages/domains, and potential biases. While our work marks an important step in medical exam question generation, broader societal impacts require careful consideration. (Semi)-Automated USMLE-MCQG systems can boost the efficiency and accuracy of medical education, but over-reliance could reduce direct student-educator interactions, and LLMs’ hallucinations might impact students’ learning and real-world practice. Caution is necessary to ensure these technologies support, not replace, medical educators. Future research should focus on these challenges to maximize positive contributions to medical education and society.

### v4

@@ 2.4 Critique
After generating all components in INIT and QA step, the LLM is asked to critique each component. The set to be critiqued is $u_{\text{critique}}=\langle c,q,a,d,r\rangle$. The LLM receives a scoring guide $G$, which includes all aspects that need to be evaluated for each component. The prompt includes this guide $G$ as well as several manually written example critiques of scored components $E^{\text{fs}}_{\text{critique}}=\langle e^{\text{fs}}_{c},e^{\text{fs}}_{q},e^{\text{fs}}_{a},e^{\text{fs}}_{d},e^{\text{fs}}_{r}\rangle$ and $\langle G,n,t,k,c,q,a,d,a_{a}\rangle$. The final output of this step is LLM critique feedback on all components, $f=\langle f_{c},f_{q},f_{a},f_{d},f_{r}\rangle$, which includes short text critiques and scores for each aspect. The aspects for scoring different components in $G$ are as follows: Context: Relevant, Concise, Coherent, Consistent, Specific, Fluent, Clueing, Completeness, and Misdirection. Question: Relevant, Clear, Concluding, Difficulty, and Clarity. Correct Answer: Relevant, Occurrence, Justification, Depth of Understanding, and Prevention of Guesswork. Distractors: Format, Length, Relation, Variation, Plausibility, Differentiation, and Common Mistakes. Chain of Thought/Reasoning: Logical Flow, Evidence-Based Reasoning, and Consideration of Options. We provide detailed explanations for each aspect of every component in Appendix Table LABEL:tab:demonstration_MCQG-SRefine, and LLM-Critique prompts in Table LABEL:prompt. The total score for each component is calculated by summing up all individual aspect scores. A sample output is provided in Appendix Table LABEL:tab:demonstration_MCQG-SRefine.
@@ 3 Experimental Design and Setup
For the medical cases utilized in the generation of USMLE questions, we used unidentified patient summaries from the PMC-Patients dataset Zhao et al. (2023). The average length of these patient summaries was ~419 words. The frequency of topics used is listed in Appendix Table 14 15.
@@
To assess the quality of the questions (RQ1), we first engaged two medical experts  [^8] Two medical students with 2+ years hospital experience. to express their preferences between the two sets (GPT-4 and GPT-4 + MCQG-SRefine) of system-generated questions based on an annotation guideline (Appendix Table LABEL:tab:human_eval_guide). The evaluators were blinded to the source of the questions, and the order of the questions was randomized for each data point. We calculated the Percentage Agreement (87.5%) and Cohen’s kappa (0.66722) for the two evaluators’ preferences, indicating substantial reliability of our human evaluation settings. Subsequently, a third human expert  [^9] One licensed physician. facilitated discussions with the initial evaluators to make the final decision for each data point, representing the final human expert preference (referred to as Expert X).
@@ 4 Results
Figure 3 demonstrates the overwhelming advantage of GPT-4 + MCQG-SRefine over GPT-4 with a 70-80% win rate in human preference about question quality (RQ1). In Figure 4, we also observe a decrease of about 80% in easy questions, a 2.25 times increase in medium questions, and a 4 times increase in hard questions with the input of these medical cases with expert-provided topics and key points. For machine-provided topics and key point cases, the proportion of easy questions decreased by 33.3%, a 2 times increase in medium questions, but there was no increase in the proportion of hard questions (RQ2). This demonstrates the effectiveness of MCQG-SRefine in increasing the difficulty of questions. This also indicates that the quality of topics and key points provided by experts is higher, so LLM can generate more difficult questions by thinking more deeply during the critique and correction steps. Further improving the quality of the machine-provided topics and key points can be a future direction for improvement.
Figure 3: The quality expert preference for the GPT-4 and the GPT-4 + MCQG-SRefine question. The data is divided into Human and Machine based on how the topic $t$ and key points $k$ were generated. We only put the final Expert X preferences here, but we provide more results in the Appendix Table 10. The percentage agreement between experts is 87.5% (Human<$t$, $k$>: 90%, Machine<$t$, $k$>: 85%). The Cohen’s kappa between experts is 0.66722 (Human<$t$, $k$>: 0.75, Machine<$t$, $k$>: 0.57), indicating substantial reliability.
@@
Figure 5 and Table 1 present the results of the LLM-as-judge evaluation. For human-provided topics and key point cases, GPT-4 + MCQG-SRefine achieved a win rate of 79.8% compared to 20.2% for GPT-4. Similarly, for machine-provided topics and key point cases, GPT-4 + MCQG-SRefine achieved a win rate of 80.1%, outperforming GPT-4’s 19.9%. Notably, GPT-4 + MCQG-SRefine consistently outperformed GPT-4 across all five evaluated components with LLM-as-judge (Rating) results rather than demonstrating an advantage in only a single aspect  [^12] The aspect-level score can be found in Appendix Table 11.. This indicates a balanced and comprehensive improvement of GPT-4 + MCQG-SRefine.
@@
Qualitative analysis One of the main issues with the questions generated by GPT-4 is that they often directly include the correct answer or too obvious relevant keywords within the context component (Appendix Table LABEL:tab:human_eval_case Case Studies 1 and 3). So the questions directly generated by GPT-4 often make the answers obvious, but GPT-4 + MCQG-SRefine can modify this information into hints for the correct answer through its critical and corrective steps, which experts consider a better way to construct questions for candidates (Case Study 4).
@@
Another finding is that the questions generated by GPT-4 + MCQG-SRefine are more concise compared to those by GPT-4 (Case Study 1). Experts pointed out that this conciseness makes them more similar to real USMLE questions. Our experiments show that GPT-4 adopted a very conservative strategy when generating context due to our emphasis on hallucination issues in the prompts. This strategy involves copying and pasting much information from raw medical case inputs to avoid generating potentially incorrect new facts. Although this does reduce the occurrence of hallucinations - our human assessment shows that 7.5% of GPT-4 problems contain factual errors - it inevitably sacrifices the typical simplicity and highly refined information presentation of USMLE problems, as well as the logical coherence of multi-hop reasoning between information in context, question, correct answer, and distractors. This is a significant stylistic difference between USMLE questions and the original medical cases. We also found that prompt engineering was ineffective in resolving this issue. We interpret this as a shortcut behavior learned by GPT-4 during aligning with human preferences stage training (e.g., RLHF Ouyang et al. (2022)) to reduce output diversity to mitigate hallucinations Kirk et al. (2023). In contrast, MCQG-SRefine maintains a high level of factual accuracy (with 5% factual errors) and further improves the quality of USMLE questions by iteratively criticizing and correcting each component of the question, as well as shifting the perspective from QG to QA to make it closer to actual exam questions. This is another major reason why GPT-4 + MCQG-SRefine outperforms GPT-4 in human evaluation. However, when MCQG-SRefine reduced the amount of useless information provided in the questions to increase quality and difficulty, omitting too much information sometimes made inferring the correct answer more challenging (Case Study 2), although this was a rare occurrence in our human evaluation (7.5%).
Figure 4: The difficulty expert evaluation for the GPT-4 generated and the GPT-4 + MCQG-SRefine questions.
5 Discussion
@@
Table 1: The win rate of GPT-4 + MCQG-SRefine and GPT-4 questions using LLM-as-Judge (Comparision).
@@ 6.1 LLMs for Generating Medical MCQs
Cheung et al. (2023) conducted the first study comparing LLMs with humans in medical exam MCQ generation. Using GPT-3.5, they generated MCQs from textbooks. GPT produced 50 MCQs in 21 minutes—10% of the time taken by humans—but human-written questions were of higher quality and effectiveness, outperforming AI in 60% of acceptable MCQs. Other studies, such as Klang et al. (2023), evaluated AI-generated MCQs using GPT-4 and found them mostly effective, though lacking in clinical context. Agarwal et al. (2023) compared GPT-3.5, Bard, and Bing in generating MCQs for a physiology course, noting that GPT produced more effective but less difficult questions. Overall, while LLMs can quickly generate MCQs, their quality and difficulty often do not meet expectations. We build on this work by qualitatively and quantitatively studying how to convert medical cases into high-quality and challenging USMLE-style MCQs.
@@ 6.2 Self-Refine using LLM Feedback
Learning from feedback helps align LLMs with desired outcomes, enhancing their ability to follow instructions through different forms of feedback, such as human preference feedback Ouyang et al. (2022), AI-generated preference feedback Lee et al. (2023); Dubois et al. (2024), or fine-grained feedback Wu et al. (2024); Lightman et al. (2023). Unlike preference and fine-grained feedback, which provide scalar values as training signals, natural language or correction feedback offers richer information Scheurer et al. (2022); Ma et al. (2023); Yao et al. (2023b); Mishra et al. (2024), making it particularly effective for self-correcting language models Welleck et al. (2022); Pan et al. (2023). Recent research has demonstrated that LLMs Achiam et al. (2023); Heng et al. (2024); Wang et al. (2024a) can self-correct their responses to meet various user requirements, such as reducing harmful content, incorporating specific keywords, diversity requirement generation, or debugging code Madaan et al. (2024); Chen et al. (2023b). This self-correction process generally involves generating a critique that identifies shortcomings in the initial response, followed by revising it based on the self-critique—an iterative process that has shown promise in enhancing LLM output quality Pan et al. (2023). Inspired by the success of these iterative self-refinement methods, we are the first to explore using this approach for generating USMLE-style MCQs. By leveraging self-feedback, we aim to create high-quality, clinically relevant questions that adhere to the rigorous standards of medical education.
@@ 6.3 LLM-as-Judge using LLM feedback
The critique capabilities of LLMs have been extensively used for the automatic evaluation of response quality, often employing models like GPT-4 Achiam et al. (2023); Liu et al. (2023); Fu et al. (2023) or critique-adjusted LLMs Ke et al. (2023); Li et al. (2023). Despite their success, these methods have demonstrated instability in certain complex task scenarios Wang et al. (2023c); Zhang et al. (2024b). LLMs have shown a high correlation with human evaluations in tasks such as summarization and story generation, effectively scoring candidate texts or comparing them based on specific evaluation aspects Chen et al. (2023a); Li et al. (2024c, b); Gu et al. (2024). For example, studies by Chiang and Lee (2023) and Kocmi and Federmann (2023) have shown that LLM evaluations produce results comparable to those of expert human evaluators in story generation and translation tasks. Similarly, research by Zheng et al. (2024) and Zhang et al. (2024a) indicates that powerful LLM reviewers, such as GPT-4, achieve over 80% consistency with human preferences in multi-turn dialogue scenarios, both in controlled and crowdsourced settings, reaching agreement levels similar to human evaluators. Further evidence of GPT-4’s effectiveness as an evaluator is demonstrated by its performance in the PROMETHEUS Kim et al. (2023, 2024) and CRITICBENCH Lan et al. (2024) benchmarks. However, most LLM-as-judge research has primarily focused on general NLP fields, with limited exploration in specialized domains like clinical NLP. This gap is largely due to challenges such as the need for domain-specific knowledge, difficulties in designing evaluation prompts that meet domain standards, and the inherent challenges of generalizing evaluation tools to specialized fields Singhal et al. (2023); Li et al. (2024d).
In the medical field Li et al. (2024c); Raju et al. (2024), LLMs-as-judges have demonstrated certain potential in areas such as diagnostic support Yao et al. (2024), medical documentation Brake and Schaaf (2024, 2024), clinical conversation Wang et al. (2023a); Li et al. (2024a), medical question answering Wang et al. (2024b); Krolik et al. (2024), medical reasoning Jeong et al. (2024), and patient education Yao et al. (2023a). To the best of our knowledge, we are the first to explore the application of LLM-as-judge in the context of USMLE-MCQ generation and evaluation. Our work addresses these challenges to create reliable, domain-specific evaluations that meet the unique requirements of medical education.
@@ 8 Limitations and Ethical Considerations
Our research presents promising advances in the generation of medical exam questions. However, several limitations and ethical considerations require further exploration to ensure the robustness, fairness, and social value of this work.
Adaptability to Other Domains and Languages: Our study focuses on medical education, specifically the USMLE format. Although the methods and frameworks show potential, their generalizability to other domains, languages, or question types remains unexplored. Future research should validate the adaptability of these methods to broader educational contexts and linguistic variations.
Evaluation Bias and Reliability of LLM-as-Judge: The LLM-as-Judge approach, where GPT-4 is used for both generating and evaluating output, carries a significant risk of self-evaluation bias. First, the evaluation and generation abilities are not equivalent. According to higher-order thinking theories  [^17] https://en.wikipedia.org/wiki/Higher-order_thinking in educational research, evaluation is a cognitively higher-dimensional task compared to task completion. Recent studies West et al. (2023) also reveal that generative AI models can excel at creating content while lacking deep understanding, which limits their reliability as evaluators. Second, while GPT-4 was selected due to its relatively smaller position bias compared to other models such as Claude3-Opus, Sonnet, Haiku, and GPT-3.5 (as shown in Appendix Table 19), it still exhibits inconsistencies in approximately 35% of cases. These inconsistencies, where GPT-4’s judgments vary based on the order of presented options (A-B vs. B-A), highlight the challenges of relying on LLMs for evaluation. Third, inconsistencies across different critique dimensions (e.g., individual scoring vs. pairwise comparison) further compound these challenges, which is aligned with recent work findings Lan et al. (2024). Addressing these limitations will require novel approaches to improve the reliability and fairness of LLM-based evaluation frameworks.
Potential for Over-Self-Critique: Our iterative self-refinement method (MCQG-SRefine) improves question quality and difficulty based on critique and correction stages. However, there is a risk of over-self-critique, where excessive iterations lead to suboptimal or overly complex outputs. While our Round-wise Analysis (Figure 6) and Appendix B highlight how question quality evolves across iterations, we observed cases where later refinements did not outperform earlier ones. This underscores the need for careful calibration of self-refinement processes to strike a balance between improvement and diminishing returns. Future work should investigate mechanisms to mitigate over-self-critique, such as introducing external evaluation checkpoints or dynamic stopping criteria.
Privacy and Ethical Considerations: The use of medical data to generate questions raises critical privacy concerns. Although our work uses publicly available clinical resources as input, and we adhere to deidentification protocols, ensuring compliance with ethical standards and safeguarding patient privacy is important. Moreover, fairness in question generation remains an open challenge. Biases present in the LLM original training data as well as input clinical notes, such as underrepresentation of certain medical conditions or demographic groups, can lead to biased outputs. Future research must prioritize fairness-aware techniques to mitigate these issues and ensure equity in medical education tools.
Impact on Medical Education and Educator-Student Dynamics: While automated USMLE-MCQG systems hold the potential to boost the efficiency, scalability, and accuracy of medical education, they may inadvertently reduce direct student-educator interactions. Medical educators are essential in providing context, clinical reasoning insights, and mentorship that automated systems cannot fully replicate. Over-reliance on automated tools could undermine these critical learning experiences. Moreover, hallucinations or inaccuracies generated by LLMs could misinform learners, posing risks to both education quality and eventual clinical practice. It is, therefore, crucial to position these tools as assistive systems that support, rather than replace, educators.
Broader Societal and Ethical Implications: The societal impacts of (semi)-automated MCQG systems extend beyond education. Ensuring that these technologies are accessible, fair, and transparent is vital to prevent exacerbating educational inequities. Moreover, the automation of question generation should be accompanied by rigorous human oversight to identify and correct potential errors. Continued collaboration with medical professionals, educators, and ethicists will be critical to addressing these challenges and ensuring the responsible deployment of AI-driven educational tools.